OpenStudy (anonymous):

The function Q(t)=Qe^-kt may be used to model radioactive decay. Q represents the quantity remaining after t years; k is the decay constant, 0.00011. How long, in years, will it take for a quantity of plutonium-240 to decay to 25% of it's original amount?

OpenStudy (anonymous):

\[Q(t)=Q _{o}e ^{-kt}\]

OpenStudy (ankit042):

q(t)= 0.25Q k =0.00011 Now you can solve for t

OpenStudy (zzr0ck3r):

Q_0 is the starting amount, (1/4)Q_0 is the ending amount \[\frac{1}{4}Q_0=Q_0e^{-0.00011t} \rightarrow \frac{1}{4}=e^{-0.00011t}\] solve for t

OpenStudy (anonymous):

I'm not sure if I did it right. Would it be approximately 12,600 years then?

OpenStudy (zzr0ck3r):

ln of both sides ln(1/4) = -0.00011t divide by -.00011 t = -ln(1/4)/0.00011

OpenStudy (zzr0ck3r):

yes -ln(1/4)/0.00011 = 12602.7

OpenStudy (anonymous):

Okay, thank you!

OpenStudy (zzr0ck3r):

im not sure what @ankit042 was doing, maybe there is another way....